OpenStudy (anonymous):

for questions 1-3, what is the solution of the system? use substitution. (will give medal and fan) 1. y=x+6 y=2x (1 point) -(6,12) -(-12,-6) -(-6,-12) -(2,4) 2. 3x+2y=7 -3x+11=y (1 point) -(6,-3) -(6,-7) -{-4,19/2} -(5,-4) 3. 2x-y=-7 4x-y=-4 (1 point ) -(-1.5,4) -(1.5,10) -(4,-1.5) -(-1.5,-2)

OpenStudy (texaschic101):

1. y = x + 6 y = 2x sub in 2x for y 2x = x + 6 2x - x = 6 x = 6 now sub 6 in for y in either of the above equations to find y y = x + 6 y = 6 + 6 y = 12 SOLUTION : (6,12) 2. 3x + 2y = 7 -3x + 11 = y sub in -3x + 11 for y in the 1st equation 3x + 2(-3x + 11) = 7 -- distribute through the parenthesis 3x - 6x + 22 = 7 -- subtract 22 from both sides 3x - 6x = 7 - 22 -- combine like terms -3x = - 15 -- divide by -3 x = 5 sub in 5 for x in either equation to find y -3x + 11 = y -3(5) + 11 = y -15 + 11 = y -4 = y SOLUTION : (5,-4) 3. 2x - y = -7 -->2x + 7 = y 4x - y = -4 now sub 2x + 7 in for y in the 2nd equation 4x - (2x + 7) = -4 -- distribute through the parenthesis 4x - 2x - 7 = -4 -- add 7 to both sides 4x - 2x = -4 + 7 -- combine like terms 2x = 3 x = 3/2 or 1 1/2 or 1.5 now sub 3/2 in for x in either of the equations 2x - y = -7 2(3/2) - y = -7 6/2 - y = -7 3 - y = -7 -- add y to each side 3 = y - 7 -- add 7 to both sides 3 + 7 = y 10 = y SOLUTION : (1.5, 10)

OpenStudy (anonymous):

Thank you so much!! youre a savior! <(^~^<)